OpenStudy (anonymous):

hellpp please Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5)

OpenStudy (anonymous):

E(-2, 2), F(5, 1) I have tried this: Y-2=(-1/7)(x-(-2) 1-2/5-(-2)=(-1/7) You found the slope: m =(1-2)/(5 +2) = -1/7

OpenStudy (anonymous):

I DONT GET IT

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

how can i get the answer

OpenStudy (anonymous):

so u dont get it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

lol oh....

OpenStudy (anonymous):

whats the formula for it?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

By slope formula: m = (1 - 2)/(5 - (-2)) ==> -1/7 Then, by y - y1 = m(x - x1) y - 1 = (-1/7)(x - 5) 7[y - 1 = -x/7 + 5/7] 7y - 7 = -x + 5 x + 7y - 7 = 5 So x + 7y = 12 I hope this helps!

OpenStudy (anonymous):

are u diling with slopes?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

https://ector.owschools.com/owsoo/studentAssignment/index This is what im doing

OpenStudy (anonymous):

u have too login

OpenStudy (anonymous):

dang never mind

OpenStudy (anonymous):

i cannot see the question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

login as 275934 and the pass is the same

OpenStudy (anonymous):

what type of pass

OpenStudy (anonymous):

275934

OpenStudy (anonymous):

password

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok now what?

OpenStudy (anonymous):

never mind lol

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

thanks anyways

OpenStudy (anonymous):

yw im trying to find someone to help me

OpenStudy (anonymous):

its ok and thanks for helping me out

OpenStudy (anonymous):

your wellcome if anything u need help on feel free to add me as a friend and good luck

OpenStudy (anonymous):

thanks and i sure will

OpenStudy (anonymous):

ok (: